# Permutation Entropy for Signal Analysis

### v2

@@ 1 Introduction
Shannon entropy was introduced in 1948 by Claude Shannon [23] to measure the uncertainty of a random variable. Specifically, given a (discrete) random variable $X$ with support $\mathcal{X}$ and PDF $p(\cdot)$, the Shannon entropy of $X$ is given by:
@@
and expresses the expected number of bits one needs to transmit to convey the outcome of $X$. Since the standard use-case for the definition of entropy is in the unit of bits, logarithms are taken to be base $2$ unless otherwise noted. Variants of Shannon entropy (e.g., Renyi entropy [6], joint entropy [10]) have been used to measure the information content of specific random variables in signals processing [15], quantum physics [27], finance [29] and many other application spaces. For a thorough treatment on the subject of entropy, we direct the reader to [10]. In this document, we focus on the so-called permutation entropy [4] (PE) as applied to time series data.
Permutation entropy as a method extracts information from time series by considering only the distribution of permutation patterns which occur in the data. In this way, PE demonstrates a concrete application of permutation pattern methods to real world data analysis problems. For example, PE has shown promise as a useful tool for signal processing with successful applications [2, 5, 7, 17, 12] ranging from biomedical to financial to mechanical. Additionally, permutation entropy has been successfully used in some Radio Frequency (RF) signal processing such as radio frequency fingerprint extraction (i.e, identifying physical radio devices from signal measurements) [11] and detecting RF transients from their fingerprint [28]. In this work we highlight the relationship between permutation patterns and RF communications modulation schemes, demonstrating the efficacy of permutation pattern-based methods on a classification problem of common RF modulations. This proof-of-concept, while a notable data science contribution in its own right, is intended to serve as a jumping off point for the use of permutation patterns in data analytics for real world RF communications use cases. This document is intended to be self-contained; the domain-specific vocabulary herein will be introduced before results are presented.
@@ 2 Time Series and Permutation Patterns
Our principal object of study is a signal. Using the definition from [20], a signal is “a function that conveys information about the behavior of a system or attributes of some phenomenon.” In this document, signals will be represented by real valued vectors (or time series) $\mathbf{x}=(x_{1},\ldots,x_{t})\in\mathbb{R}^{t}$, where $\{x_{i}\}_{i=1}^{t}$ are typically data points sampled at equal time intervals from some system or sensor [^1] Notationally, many of our definitions are greatly simplified by taking the starting index of $1$, while much of the real world data we examine comes 00-indexed. For symbolic definitions, we use $1$-indexing, but specific data analysis is displayed using 00-indexing. The indexing used will change as needed, but will be made clear from context whenever possible.. Frequently, the (signed) magnitude of the values of $\mathbf{x}$ (i.e., amplitudes) are normalized, presenting an additional preprocessing step to make them more comparable when they are structurally similar but have very different scales. This normalization is often non-trivial due to the unknown global characteristics of the signals. PE is a function of the relative positions of the entries of $\mathbf{x}$ to one another, and is thus amplitude agnostic. Thus, PE is a function of relative patterns in a signal without suffering inefficiencies of scale or amplitdudinal noise. We make the notion of a signal pattern precise in the following definition:
@@
Note that $H_{n,\tau}(\cdot)$ is maximized when the permutation distribution is uniform [8] with a value of $\log n!$. Entropy is typically a number between 00 and $1$; to this end, we introduce the normalized $n$,$\tau$ permutation entropy (NPE):
@@
In practice, the choice of $n$ and $\tau$ is not trivial [18, 25]. One remedy is to instead use Multi-Scale Permutation Entropy (MSPE) in which multiple combinations of $\tau$ and $n$ are used to compute a matrix of permutation entropy values. For example if we use the same signal as in Fig. 2(a) but instead of a single $n=3$ and $\tau=2$ we choose $n\in[2,3]$ and $\tau\in[1,2]$ we get the MSPE matrix
@@
where the rows are for each dimension $n$ and the columns are for each delay $\tau$. This matrix now holds additional information that would have been lost through the choice of a single dimension and delay. For the interested reader, an early application of multiscale entropy can be found in [9] which deals with physiological time series (e.g., heart rate). An early application of MSPE can be found in [3] with the same biomedical use case. This paper includes the first application of multi-scale permutation entropy to the classification of RF signal modulations, which will be explained in Section 3.
@@ 3 RF signal processing background
Radio frequency (RF) waves model the oscillation of various media including electrical, electromagnetic, or mechanical systems in the range of $30$kHz to about $300$GHz [21]. Familiar applications of RF signals include AM/FM radio, television, bluetooth, and internet communications. RF digital communications are generated by modulating a binary sequence into a wave-form that can be transmitted and received (see Fig. 3 for an example) to relay the desired information. We make this specification because this use-case (e.g., sequences of bits) is pervasive, but also because entropy is a measurement of the communication of information, and thus entropic quantification is a natural method for studying RF signals. We will not go into details about specific frequencies or transmission methods– an introduction on RF signals can be found here [26]. Rather, we provide a high level explanation of the process for sending binary signals via RF signals to inform our experiments.
@@
BPSK modulators take as input a binary sequence and a carrier signal. Instances of the carrier signal in two different phases are concatenated, with one phase representing $1$ bits and another representing 00 bits. The modulated signal is transmitted, and the received signal (which is a close approximation of the transmitted signal) is turned into a signal which is positive during periods meant to represent a $1$, and negative otherwise. The recovered signal is translated back to a binary string accordingly. For details about the implementation of BPSK, see [24]. We have included a simple example in Figure 3.
@@ 4 Methods
Here, $\ell$ is the largest integer such that $s_{\ell}+(k-1)\leq t$. Informally, we window $\mathbf{x}$ and compute the permutation entropy for each window. In this setup $n$, $\tau$, $k$, and $\alpha$ are tuneable parameters. The parameters $n$ and $\tau$ are frequently selected according to heuristics or are informed by domain knowledge. A collection of automatic parameter selection tools for selecting $n$ and $\tau$ can be found in [18]. Suitable choices for $k$ and $\alpha$ vary based on use cases and available computation. In general, if $k$ is too small then the distribution of permutation patterns is not very informative, and if $\alpha$ is too small then the sequence of permutation entropies is coarse and can vary greatly from term to term. However, the computation of the permutation distribution scales with $k$ and the number of windows scales with $\alpha$, so taking both $k$ and $\alpha$ large can be computationally expensive.
@@ 5 Data and Results
We introduce in this section the dataset used to evaluate our MSPE approach, give futher detail about our evaluation method, and present the results of a comparison study where modest Convolutional Neural Networks (CNNs) were trained to classify modulation schemes when given MPSE, spectrogram, or raw waveform data as input. The data we work with is the Panoradio data set introduced in [22] (Scholl) and sourced at [1]. [^4] The purpose of this article is to show the connection between permutation patterns and signal analysis. To this end, we present only a high level overview of the Panoradio data to demonstrate efficacy of permutation entropy on a classification task. However, this data is thoroughly documented, and the interested reader is encouraged to explore the source material [22] [1]. This dataset is an extension of a more standard modulation classification dataset [19] which further incorporates different modes of communication and baud rates [^5] Baud rate: modulation rate in units of symbols-per-second. Specifically, there are 18 different modulation modes in the data set that are shown in Table 1 in the appendix with both the modulation method and the baud rate if applicable. Additionally, Fig. 7 of the appendix provides example signals from each of the 18 modulation methods. These modes cover many of the standard modulation methods used in modern RF signal processing. Plain text was encoded for digital modes, speech and music were used for analog modes, and black and white images for fax data. To make this simulated dataset more realistic there is additive white Gaussian noise at different levels, non-coherent reception and receiver frequency mismatch, and random phase and frequency offsets.
@@
The CNNs we use have two convolution layers, each followed by maximum pooling, and one fully connected layer. We trained each CNN for 30 epochs with a batch size of 10. The CNN trained on the raw signal, MSPE, and spectrograms had a total of 37.0K, 12.6K, and 36.5K parameters, respectively. In comparison to the original Panoradio dataset [22] with approximately 1.4 million parameters and 17 layers, these neural networks are significantly smaller and quicker to train.
@@ 6 Discussion
While the modulation classification performance serves as a proof-of-concept that the distribution of permutation patterns which occur within time series signal data captures critical information about RF signals (with simple MSPE models beating large CNN and ResNet models on raw data), modulation dependent methods face limitations in practice. For example, the 18 modulations in the Panoradio data set are in no way comprehensive; many use cases come with bespoke modulation schemes and adversarial methods can be used to confound modulation detection schema [16]  [14]. However, minor signal perturbations should not drastically affect the distribution of permutation patterns within a signal. A qualitative analysis of which properties of a signal that MSPE methods capture is the first step towards modulation agnostic signal information analysis which is an active and ongoing research area.
@@ 7 Future Permutation Patterns Work
Combinatorially, we find that the converse question of time series which are realizable given a distribution of permutation patterns has connections to covering, packing, and universal sequences [13]. To this end, we ask:

### v3

@@ 1 Introduction
Shannon entropy was introduced in 1948 by Claude Shannon Shannon (2001) to measure the uncertainty of a random variable. Specifically, given a (discrete) random variable $X$ with support $\mathcal{X}$ and PDF $p(\cdot)$, the Shannon entropy of $X$ is given by:
@@
and expresses the expected number of bits one needs to transmit to convey the outcome of $X$. Since the standard use-case for the definition of entropy is in the unit of bits, logarithms are taken to be base $2$ unless otherwise noted. Variants of Shannon entropy (e.g., Renyi entropy Bromiley et al. (2004), joint entropy Cover (1999)) have been used to measure the information content of specific random variables in signals processing Hughes (1993), quantum physics Van Herstraeten and Cerf (2021), finance Zhou et al. (2013) and many other application spaces. For a thorough treatment on the subject of entropy, we direct the reader to Cover (1999). In this document, we focus on the so-called permutation entropy Bandt and Pompe (2002) (PE) as applied to time series data.
Permutation entropy as a method extracts information from time series by considering only the distribution of permutation patterns which occur in the data. In this way, PE demonstrates a concrete application of permutation pattern methods to real world data analysis problems. For example, PE has shown promise as a useful tool for signal processing with successful applications Amigó et al. (2012); Bariviera et al. (2018); Cao et al. (2004); Morabito et al. (2012); Garland et al. (2018) ranging from biomedical to financial to mechanical. Additionally, permutation entropy has been successfully used in some Radio Frequency (RF) signal processing such as radio frequency fingerprint extraction (i.e, identifying physical radio devices from signal measurements) Deng et al. (2017) and detecting RF transients from their fingerprint Yuan et al. (2015). In this work we highlight the relationship between permutation patterns and RF communications modulation schemes, demonstrating the efficacy of permutation pattern-based methods on a classification problem of common RF modulations. This proof-of-concept, while a notable data science contribution in its own right, is intended to serve as a jumping off point for the use of permutation patterns in data analytics for real world RF communications use cases. This document is intended to be self-contained; the domain-specific vocabulary herein will be introduced before results are presented.
@@ 2 Time Series and Permutation Patterns
Our principal object of study is a signal. Using the definition from Priemer (1991), a signal is “a function that conveys information about the behavior of a system or attributes of some phenomenon.” In this document, signals will be represented by real valued vectors (or time series) $\mathbf{x}=(x_{1},\ldots,x_{t})\in\mathbb{R}^{t}$, where $\{x_{i}\}_{i=1}^{t}$ are typically data points sampled at equal time intervals from some system or sensor [^1] Notationally, many of our definitions are greatly simplified by taking the starting index of $1$, while much of the real world data we examine comes 00-indexed. For symbolic definitions, we use $1$-indexing, but specific data analysis is displayed using 00-indexing. The indexing used will change as needed, but will be made clear from context whenever possible.. Frequently, the (signed) magnitude of the values of $\mathbf{x}$ (i.e., amplitudes) are normalized, presenting an additional preprocessing step to make them more comparable when they are structurally similar but have very different scales. This normalization is often non-trivial due to the unknown global characteristics of the signals. PE is a function of the relative positions of the entries of $\mathbf{x}$ to one another, and is thus amplitude agnostic. Thus, PE is a function of relative patterns in a signal without suffering inefficiencies of scale or amplitdudinal noise. We make the notion of a signal pattern precise in the following definition:
@@
Note that $H_{n,\tau}(\cdot)$ is maximized when the permutation distribution is uniform Conrad (2004) with a value of $\log n!$. Entropy is typically a number between 00 and $1$; to this end, we introduce the normalized $n$,$\tau$ permutation entropy (NPE):
@@
In practice, the choice of $n$ and $\tau$ is not trivial Myers and Khasawneh (2020); Staniek and Lehnertz (2007). One remedy is to instead use Multi-Scale Permutation Entropy (MSPE) in which multiple combinations of $\tau$ and $n$ are used to compute a matrix of permutation entropy values. For example if we use the same signal as in Fig. 2(a) but instead of a single $n=3$ and $\tau=2$ we choose $n\in[2,3]$ and $\tau\in[1,2]$ we get the MSPE matrix
@@
where the rows are for each dimension $n$ and the columns are for each delay $\tau$. This matrix now holds additional information that would have been lost through the choice of a single dimension and delay. For the interested reader, an early application of multiscale entropy can be found in Costa et al. (2002) which deals with physiological time series (e.g., heart rate). An early application of MSPE can be found in Aziz and Arif (2005) with the same biomedical use case. This paper includes the first application of multi-scale permutation entropy to the classification of RF signal modulations, which will be explained in Section 3.
@@ 3 RF signal processing background
Radio frequency (RF) waves model the oscillation of various media including electrical, electromagnetic, or mechanical systems in the range of $30$kHz to about $300$GHz Scarpati (2021). Familiar applications of RF signals include AM/FM radio, television, bluetooth, and internet communications. RF digital communications are generated by modulating a binary sequence into a wave-form that can be transmitted and received (see Fig. 3 for an example) to relay the desired information. We make this specification because this use-case (e.g., sequences of bits) is pervasive, but also because entropy is a measurement of the communication of information, and thus entropic quantification is a natural method for studying RF signals. We will not go into details about specific frequencies or transmission methods– an introduction on RF signals can be found here Tse and Viswanath (2005). Rather, we provide a high level explanation of the process for sending binary signals via RF signals to inform our experiments.
@@
BPSK modulators take as input a binary sequence and a carrier signal. Instances of the carrier signal in two different phases are concatenated, with one phase representing $1$ bits and another representing 00 bits. The modulated signal is transmitted, and the received signal (which is a close approximation of the transmitted signal) is turned into a signal which is positive during periods meant to represent a $1$, and negative otherwise. The recovered signal is translated back to a binary string accordingly. For details about the implementation of BPSK, see Sklar (2021). We have included a simple example in Figure 3.
@@ 4 Methods
Here, $\ell$ is the largest integer such that $s_{\ell}+(k-1)\leq t$. Informally, we window $\mathbf{x}$ and compute the permutation entropy for each window. In this setup $n$, $\tau$, $k$, and $\alpha$ are tuneable parameters. The parameters $n$ and $\tau$ are frequently selected according to heuristics or are informed by domain knowledge. A collection of automatic parameter selection tools for selecting $n$ and $\tau$ can be found in Myers and Khasawneh (2020). Suitable choices for $k$ and $\alpha$ vary based on use cases and available computation. In general, if $k$ is too small then the distribution of permutation patterns is not very informative, and if $\alpha$ is too small then the sequence of permutation entropies is coarse and can vary greatly from term to term. However, the computation of the permutation distribution scales with $k$ and the number of windows scales with $\alpha$, so taking both $k$ and $\alpha$ large can be computationally expensive.
@@ 5 Data and Results
We introduce in this section the dataset used to evaluate our MSPE approach, give futher detail about our evaluation method, and present the results of a comparison study where modest Convolutional Neural Networks (CNNs) were trained to classify modulation schemes when given MPSE, spectrogram, or raw waveform data as input. The data we work with is the Panoradio data set introduced in Scholl (2019b) and sourced at Scholl (2019a). [^4] The purpose of this article is to show the connection between permutation patterns and signal analysis. To this end, we present only a high level overview of the Panoradio data to demonstrate efficacy of permutation entropy on a classification task. However, this data is thoroughly documented, and the interested reader is encouraged to explore the source material Scholl (2019b) Scholl (2019a). This dataset is an extension of a more standard modulation classification dataset O’Shea et al. (2018) which further incorporates different modes of communication and baud rates [^5] Baud rate: modulation rate in units of symbols-per-second. Specifically, there are 18 different modulation modes in the data set that are shown in Table 1 in the appendix with both the modulation method and the baud rate if applicable. Additionally, Fig. 7 of the appendix provides example signals from each of the 18 modulation methods. These modes cover many of the standard modulation methods used in modern RF signal processing. Plain text was encoded for digital modes, speech and music were used for analog modes, and black and white images for fax data. To make this simulated dataset more realistic there is additive white Gaussian noise at different levels, non-coherent reception and receiver frequency mismatch, and random phase and frequency offsets.
@@
The CNNs we use have two convolution layers, each followed by maximum pooling, and one fully connected layer. We trained each CNN for 30 epochs with a batch size of 10. The CNN trained on the raw signal, MSPE, and spectrograms had a total of 37.0K, 12.6K, and 36.5K parameters, respectively. In comparison to the original Panoradio dataset Scholl (2019b) with approximately 1.4 million parameters and 17 layers, these neural networks are significantly smaller and quicker to train.
@@ 6 Discussion
While the modulation classification performance serves as a proof-of-concept that the distribution of permutation patterns which occur within time series signal data captures critical information about RF signals (with simple MSPE models beating large CNN and ResNet models on raw data), modulation dependent methods face limitations in practice. For example, the 18 modulations in the Panoradio data set are in no way comprehensive; many use cases come with bespoke modulation schemes and adversarial methods can be used to confound modulation detection schema Lin et al. (2020)  Hameed et al. (2020). However, minor signal perturbations should not drastically affect the distribution of permutation patterns within a signal. A qualitative analysis of which properties of a signal that MSPE methods capture is the first step towards modulation agnostic signal information analysis which is an active and ongoing research area.
@@ 7 Future Permutation Patterns Work
Combinatorially, we find that the converse question of time series which are realizable given a distribution of permutation patterns has connections to covering, packing, and universal sequences Godbole et al. (2018). To this end, we ask: